OpenStudy (anonymous):

Find a polynomial of degree n that has the given zeros. (There are many correct answers.) zeroes: x=-1,5,7 degree: n=3

OpenStudy (anonymous):

We use the fact that if a is a zero then x-a is a factor of the polynomial so product of all the factors will give us the required polynomial

OpenStudy (anonymous):

(x+1)(x-5)(x-7) = 0 is the simplest equation expand this to get polynomial

OpenStudy (anonymous):

so polynomial will be k[(x+1)(x-5)(x-7)] so now find the product of the three binomials in the bracket n you get yr answer

OpenStudy (anonymous):

my problem was the expansion. i knew what you had to do. i just dont know how to expand correctly.

OpenStudy (anonymous):

okay ill help you ☺

OpenStudy (anonymous):

(x+1)(x-5)(x-7) first we multiply (x+1)(x-5) = x(x-5) +1(x-5) = x^2 - 5x + x - 5 = x^2 - 4x -5

OpenStudy (anonymous):

now we multiply the above result by the third factor (x^2 - 4x - 5)(x-7) = x^2(x-7) -4x(x-7) -5(x-7) = x^3 - 7x^2 -4x^2 + 28x -5x + 35 = x^3 - 11x^2 + 23x + 35 Hope it is clear now....☺☻

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

u r welcome..